# Data-Adaptive Identification of Subpopulations Vulnerable to Chemical Exposures using Stochastic Interventions

### v1

@@ 7.8 Pooled Oracle Estimates
For each region and it’s complement, we provide the pooled oracle estimates. Here, $v$ denotes the regions for lower age and $v^{c}$ for higher age. Below is Table 2 of these pooled results.
@@
This indicates that the average impacts of shifting 3,3’,4,4’,5-pcnb by 1 standard deviation, -52.8, leads to an increase in telomere length of 0.488 in the lower age group, demarcated on average as centered age of 14, and of 0.316 in the higher age group.
We can compare the two groups based on $\Psi$ and the standard error using a two-sample z-test. For these results, the z-score is 5.06 resulting in a corresponding p-value p $<0.001$, indicating a significant difference between the two groups.
@@ 8 Conclusion
Our study presents a novel methodology, implemented in the open-source package, EffectXshift, found on github, to identify subpopulations that are differentially affected by chemical exposures and to estimate the effects of interventions within these subpopulations. Methodologically, our approach addresses several limitations of traditional causal inference methods by providing a data-driven, assumption-lean framework that leverages machine learning for both the estimation of exposure effects and the identification of effect modifiers. The use of stochastic shift interventions within our framework allows for more realistic and flexible modeling of exposure distributions, which is particularly relevant in environmental health studies, where exposures are often not binary or fixed.
@@
Despite the strengths of our approach, there are some limitations. The accuracy of our estimates depends on correct estimation of the clever covariate (ratio of densities), which can be difficult and the reason we offer an alternative approach via the classification approach. However, more research is necessary to ensure this reparameterization of the density estimation is sound in practice, although we show very little difference in results; our DGP is relatively simple. Additionally, while our method can handle complex and high-dimensional data, the computational cost may be significant for very large datasets. Future research should focus on further optimizing the computational efficiency of the EffectXshift package and exploring its applicability to other types of environmental exposures and outcomes.
In conclusion, the EffectXshift package provides a powerful tool for environmental health researchers to identify and estimate the differential impacts of chemical exposures between subpopulations. By facilitating the discovery of vulnerable groups and informing targeted interventions, this methodology contributes to the broader goal of reducing health disparities and improving public health outcomes.

### v2

@@ 6.4 Pooled Oracle Estimates
For each region and its complement, we provide the pooled oracle estimates. Here, $v$ denotes the region for the lower age group and $v^{c}$ for the higher age group. Below is Table 2 of these pooled results.
@@
These estimates indicate that shifting 3,3’,4,4’,5-pcnb by one standard deviation (approximately -52.8) leads to an estimated increase in telomere length of about 0.488 in the lower age group ($v$) and about 0.316 in the higher age group ($v^{c}$).
To formally assess whether the difference between these two subgroups is statistically significant, we can compare the estimates and their variances directly. Letting $\hat{\psi}_{\mathcal{V}}=\hat{\Psi}_{v}-\hat{\Psi}_{v^{c}}=0.4879-0.3156=0.1723$, and assuming negligible covariance between the estimates, the variance of the difference is $\widehat{\mathrm{Var}}(\hat{\psi}_{\mathcal{V}})=0.0001027+0.0010592=0.0011619$. Thus, $\text{SE}(\hat{\psi}_{\mathcal{V}})=\sqrt{0.0011619}\approx 0.0341$.
The corresponding z-statistic is:
$Z=\frac{0.1723}{0.0341}\approx 5.06,$
which yields a p-value well below 0.001. This confirms that the difference in the effect estimates between the two subpopulations is highly statistically significant.
@@ 7 Conclusion
Our study presents a novel methodology, implemented in the open-source package, EffectXshift, found on github, to identify subpopulations that are differentially affected by exposures and to estimate the effects of interventions within these subpopulations. Methodologically, our approach addresses several limitations of traditional causal inference methods by providing a data-driven, assumption-lean framework that leverages machine learning for both the estimation of exposure effects and the identification of effect modifiers. The use of stochastic shift interventions within our framework allows for more realistic and flexible modeling of exposure distributions, which is particularly relevant in environmental health studies, where exposures are often not binary or fixed.
@@
Despite the strengths of our approach, there are some limitations. The accuracy of our estimates depends on the correct estimation of the clever covariate (the ratio of densities). Estimating this ratio can be challenging, which is why we offer an alternative approach via classification. However, more research is necessary to ensure this reparameterization of the density estimation is sound in practice, although we show very little difference in results; our DGP is relatively simple. Additionally, while our method can handle complex and high-dimensional data, the computational cost may be significant for very large datasets. Future research should focus on further optimizing the computational efficiency of the EffectXshift package and exploring its applicability to other types of environmental exposures and outcomes.
In conclusion, the EffectXshift package provides a powerful tool for environmental health researchers to identify and estimate the differential impacts of exposures between subpopulations. By facilitating the discovery of vulnerable groups and informing targeted interventions, this methodology contributes to the broader goal of reducing health disparities and improving public health outcomes.